# SoK: A Classification for AI-driven Personalized Privacy Assistants

### v2

@@ 3.2.1. Search Strategy
Based on our RQs and previous preliminary searches when designing the SLR Protocol, we identified a list of nine relevant keywords, i.e., privacy, data protection, assistant, agent, artificial intelligence, machine learning, intelligent, automatic, and personalized. These keywords were used to construct the following search query: {spverbatim} (privacy OR ”data protection”) AND (assistant* OR agent*) AND (”artificial intelligence” OR ”machine*learning” OR intelligent OR automat* OR personali*ed) As such, the search query targets papers working on three joint topics: 1) privacy (or data protection), using either 2) an assistant or an agent, and leveraging 3) artificial intelligence or personalization.
@@
Four scientific databases were selected, i.e., Scopus, Web of Science, IEEE Xplore, and ACM Digital Library, due to their high relevance to the areas of computer science and engineering, comprising the vast majority of published research in the field. We also specified inclusion and exclusion criteria (see Table 1) used during the screening of publications retrieved from the databases. Before starting the search process, two authors piloted the searches on all databases and ran a calibration exercise to verify the consistency of the inclusion criteria. For that, the authors independently screened 10% of the results and discussed their decisions. The conflicts were all discussed and solved, sometimes with the help of a third author. This process was repeated a second time, screening another 10% of the papers at a point that the authors fully agreed with the consistency of the selection process.
@@
Table 1. Criteria for the inclusion and exclusion of studies.
3.2.2. Selection Process
@@
The first data extraction phase, consisting of a full reading of each of the 33 papers, was performed over weeks 4 to 7 (included) in 2024. Fifteen papers were excluded after full reading for different reasons: they were duplicates (i.e., same work published in different venues); they did not provide any technical solution; the automated decisions were not personalized to an end user; AI was not used for automating decisions; or they are of poor scientific quality (see our criteria in Table 1); one paper was not available for download, we could not access it even after reaching out the authors.
@@
We then proceeded to several snowballing phases (Wohlin, ), during which we checked the abstracts of all seemingly relevant [^2] We only assessed papers cited in relevant sections such as the related work. papers cited (backward snowballing), and screened citing papers (forward snowballing). The snowballing process lasted from week 8 to week 19 of 2024 and resulted in 21 additional papers after exclusion, for a total of 39 papers (33-15+21).
3.2.3. Data Extraction and Analysis

### v4

@@ III-B1 Search Strategy
Based on our RQs and previous preliminary searches when designing the SLR Protocol, we identified a list of nine relevant keywords, i.e., privacy, data protection, assistant, agent, artificial intelligence, machine learning, intelligent, automatic, and personalized. These keywords were used to construct the search query in Listing 1. As such, the search query targets papers working on three joint topics: 1) privacy (or data protection), using either 2) an assistant or an agent, and leveraging 3) artificial intelligence or personalization.
Search Query = {
(privacy OR "data protection") AND
(assistant* OR agent*) AND ("artificial
intelligence" OR "machine*learning"
OR intelligent OR automat* OR personali*ed)
Listing 1: Composition of Search Query for Literature Search.
@@
Four scientific databases were selected, i.e., Scopus, Web of Science, IEEE Xplore, and ACM Digital Library, due to their high relevance to the areas of computer science and engineering, comprising the vast majority of published research in the field. We also specified inclusion and exclusion criteria (see Table I) used during the screening of publications retrieved from the databases. Marky et al. [17] is a good illustration of a relevant paper not meeting our selection criteria. In spite of providing a technical solution to automate privacy decisions (explicitly called a PPA), the paper does not use AI and hence is not included in our list of papers. Bollinger et al. [18] provides another example of an excluded yet relevant paper. The paper provides a technical solution for automating privacy decisions and uses AI, but does not personalize the decisions.
TABLE I: Criteria for the inclusion and exclusion of studies.
@@
Before starting the search process, two authors piloted the searches on all databases and ran a calibration exercise to verify the consistency of the inclusion criteria. For that, the authors independently screened 10% of the results and discussed their decisions. The conflicts were all discussed and solved, sometimes with the help of the third author. This process was repeated a second time, screening another 10% of the papers at a point that the authors fully agreed with the consistency of the selection process.
III-B2 Data management
To manage the screening process, we exported search results from each database and imported them to the RAYYAN software (https://rayyan.ai/), allowing two reviewers to independently select papers (i.e., double-blinded) and to manage conflicts by a third reviewer. Duplicated publications were also removed using RAYYAN during the selection process. Bibliographies of final results were exported to Zotero (for citing and sharing research).
III-B3 Selection Process
@@
The first data extraction phase, consisting of a full reading of each of the 33 papers, was performed over weeks 4 to 7 (included) in 2024. Fifteen papers were excluded after full reading for different reasons: they were duplicates (i.e., same work published in different venues); they did not provide any technical solution; the automated decisions were not personalized to an end user; AI was not used for automating decisions; or they are of poor scientific quality (see our criteria in Table I); one paper was not available for download, we could not access it even after reaching out the authors.
@@
We then proceeded to several snowballing phases [19], during which we checked the abstracts of all seemingly relevant [^3] We only assessed papers cited in relevant sections, e.g., related work. papers cited (backward snowballing), and screened citing papers (forward snowballing). The snowballing process lasted from week 8 to week 19 of 2024 and resulted in 21 additional papers after exclusion, for a total of 39 papers (33-15+21). We performed an update of our results in February 2025, repeating the whole process in a 2nd round of searches. This update led us to include one additional paper through the database searches and another one through the snowballing process, for a total of 41 papers.
III-B4 Data Extraction and Analysis